# GW/DT invariants and 5D BPS indices for strips from topological recursion

### v1

@@ Abstract.
Topological string theory partition function gives rise to Gromov-Witten invariants, Donaldson-Thomas invariants and 5D BPS indices. Via the remodeling conjecture connecting Topological Recursion and topological string theory for toric Calabi-Yau 3-folds, a more direct connection is studied for the subclass of strip geometries. In doing so, new developments in the theory of topological recursion are applied as its extension to Logarithmic Topological Recursion (Log-TR) and the universal xxyy duality. Through these techniques, our main result in this paper is a direct derivation of all free energies from topological recursion for general strip geometries. In analyzing the expression of free energy, we shed some light on the meaning and the influence of the xxyy duality in topological string theory and its interconnection to GW and DT invariants as well as the 5D BPS index.
@@ 1. Introduction and summary
In this work, our main focus lies on understanding the connection among various kinds of invariants, motivated by enumerative geometry and physics, associated with Calabi-Yau threefolds (CY). Namely, the Gromov-Witten (GW) invariants arising as enumerative invariants in the context of counting stable maps from a Riemann surface $\Sigma_{g}$ with genus $g$ into the CY ($\mathcal{Y}$), in the class $\beta\in H_{2}(\mathcal{Y},\mathbb{Z})$, on the one hand. Through the “remodeling conjecture” [BKMnP09], (several aspects of which have been firmly established as mathematically proven already [EO15, FLZ15, FLZ20]) the GW invariants can be associated with the B-model geometry of the mirror $\mathcal{X}$, as periods/variation of Hodge-structure. On the other hand, we consider the Donaldson-Thomas (DT) invariants for a CY 3-fold $\mathcal{Y}$, associated with the counting of coherent sheaves, with stability determined by polarization [JS12, Beh09]. As we will see, they are also closely related to the counting of certain BPS indices in five-dimensional supersymmetric gauge theories, engineered by CY 3-folds [NO16],[Nek17],[BLR19]. In the following, we will denote as $\mathcal{X}$ the mirror of the CY 3-fold $\mathcal{Y}$. In fact, the same counting problem through mirror symmetry can be formulated as the counting of stable special Lagrangian A-branes in the Calabi-Yau threefold ($\mathcal{X}$).
@@
The DT invariants correspond to the counting of BPS states in the 5D $\mathcal{N}=1$ theory obtained by compactifying M-theory on a CY threefold. Starting from M-theory on $\mathcal{Y}\times S^{1}\times\mathbb{R}^{4}$, one arrives at the 5D theory, defined by the geometry of the CY as $T^{\textrm{5D}}[\mathcal{Y}]$ on ${\mathbb{R}}^{4}\times S^{1}$. Now consider M5-branes wrapping $\mathcal{C}_{4}\times S^{1}\times\mathbb{R}$ and M2-branes wrapping $\mathcal{C}_{2}\times{\mathrm{pt}}\times\mathbb{R}$, where $\mathcal{C}_{i}$ are holomorphic submanifolds in $\mathcal{Y}$ (ii corresponds to real dimension) and the second component and the third component indicate the submanifolds that they wrap inside $S^{1}$ and $\mathbb{R}^{4}$. Reducing on $S^{1}$ one obtains a bound state of D4-D2-D0 branes in type IIA string theory corresponding to an object of the derived (bounded) category of coherent sheaves $D^{b}Coh(\mathcal{Y})$ [DVV06, KS00, Kon95b, NO16]. At the same time in $T^{\textrm{5D}}[\mathcal{Y}]$, the corresponding BPS objects are “monopole strings” wrapping $S^{1}\times\mathbb{R}$ and “instanton particles” wrapping $\mathbb{R}$. This relates the computation of the BPS index for this theory, to the computation of the “rank zero DT invariants” for the Calabi-Yau threefold $\mathcal{Y}$. In this work, for a specific class of geometries, we will provide these 5D BPS indices $\Omega$, starting from the GW partition function.
@@
The type IIA string theory setup for the count of the $\Omega$ is related to the type IIA setup of the Gopakumar-Vafa (GV) through “9-11” flip [DVV06]. Starting from the background $\mathcal{Y}\times S^{1}\times\mathbb{R}^{4}$, one deforms the $\mathbb{R}^{4}$ as a Taub-Nut as $\mathbb{R}^{3}\times S^{1}_{\mathrm{TN}}$. Reducing on the circle $S^{1}_{\mathrm{TN}}$ now, one obtains the relevant type IIA frame where D2 and D0 branes lift to M2 branes and Kaluza-Klein momenta in M-theory. It is possible to include D4 branes which lift to M5 branes. However, for the class of CY threefolds $\mathcal{Y}$ that we will consider, there will be no compact 4-cycle, so we omit their description here.
@@
The GV invariants are intimately related to the GW invariants. This is well-known from both physics and mathematics perspectives [Kon95a, Pan03]. The GW invariants however appear naturally in the context of A-model where one counts stable maps $\phi:\Sigma_{g}\to\mathcal{Y}$. Then the path integral in the A-model localizes to the holomorphic maps which correspond to closed algebraic curves in $\mathcal{Y}$. We will return to their geometric meaning in a short while, after describing the class of the CY threefolds that in which our interest lies.
@@
We consider toric CY threefolds $\mathcal{Y}$ which in the A-model side appear as a symplectic quotient $\mathcal{Y}=\mathbb{C}^{k+3}//U(1)^{k}$. The moment map equations give rise to a toric diagram which define $\mathcal{Y}$. The mirror B-model description was provided in [HV00], which was $\mathcal{X}=\{A(e^{x},e^{y})=uv\}\subset\mathbb{C}^{4}$. The mirror Calabi-Yau $\mathcal{X}$ then can be considered as a conic fibration on a Riemann surface defined through $A(e^{x},e^{y})=0$. In the particular class of examples that we consider, we focus on the “strip geometries” which belong to the class of “toric trees” [FJ05, PS19, IKP06]. They do not have any compact 4-cycle and the “mirror curve” is given as follows [PS19]
@@
One way to relate the problem of counting 5D indices $\Omega$ (which are related to the DT invariants) is making the problem slightly more complicated through introducing a toric brane, which we will refer to as Aganagic-Vafa (AV) brane [AV00, AKV02, BKMnP09]. Consider an M5 brane wrapping $L\times S^{1}\times\mathbb{R}^{2}$, where $L$ corresponds to a special Lagrangian A-brane with topology $\mathbb{C}\times S^{1}$. It was argued in [AV00] that the complexified moduli space of this toric brane is the mirror curve corresponding to the mirror CY 3-fold $\mathcal{X}$. This is at the heart of the remodeling conjecture proposed in [BKMnP09] connecting the B-model computation to topological recursion. Our computations in this paper clarify certain aspects of it, in the context of strip geometries.
However the presence of the AV brane gives rise to a new BPS sector corresponding to a 3D $\mathcal{N}=2$ theory engineered by the codimension two defect coupled to the 5D theory, giving rise to the 3D-5D BPS sector. These were partly understood in [BLR19, BLR21, BLR20]. Generalizing the works of [GMN13], it was shown how to compute the DT invariants of $\mathcal{X}$ starting from $(\mathcal{Y},L_{\mathrm{AV}})$ through the procedure of “nonabelianization”. This later led to a geometric definition of DT invariants corresponding to objects in the Fukaya category $D^{b}Fuk(\mathcal{X})$ which are compactly supported special Lagrangian A-branes, [BLR23], providing evidence for a proposal of Joyce [Joy02]. An alternative perspective emerged in [JL25] and more will be clarified in an upcoming work [BIJ], whereas the computations here will be extended for the case of open GW invariants and connections to open DT invariants will be clarified in [BH].
As was said, we focus on toric tree Calabi-Yau (CY) threefolds that do not have compact 4-cycles given by a set of $\mathbb{P}^{1}$’s which have resolutions as either $\mathcal{O}\oplus\mathcal{O}(-2)\to\mathbb{P}^{1}$ or $\mathcal{O}(-1)\oplus\mathcal{O}(-1)\to\mathbb{P}^{1}$. The topological string partition functions for such geometries are known [IKP06, OSY11]. Let us recall them here:
@@
where we put $s_{i_{k}}=-1$ if the $\mathbb{P}^{1}$ was resolved by $\mathcal{O}(-1,-1)$ and $s_{i_{k}}=+1$ if the resolution is $\mathcal{O}(0,-2)$. Here $\chi$ is the Euler characteristic of the CY and $Q_{i}$ are the complex structure parameters. Looking at the partition function itself, we can write down the toric tree unambiguously, because the topological string partition function can be built up just from the gluing data. Expanding the partition function in terms of the free energy by $Z=\exp\sum_{g}\hbar^{2g-2}F_{g}$ with $q=e^{\hbar}$ yields for $g\geq 2$
@@
for some $k\in\{1,...,\chi-1\}$ depending on the structure of the strip geometry. The 5D BPS invariants depend on the charge $\Gamma(\mathcal{E})={\textrm{ch}}(\mathcal{E})\sqrt{\textrm{Td}(\mathcal{X})}$, for some coherent sheaf $\mathcal{E}$, $\mathcal{X}$ being the CY 3-fold in the B-model side. From this one can read off the charge vectors $\gamma$ which define the D4-D2-D0 content. Using our proposal, we read off $\Omega(\gamma)$ from (1.3) (in these examples D4-brane charge is zero).
As an example, consider the five punctured sphere obtained from gluing two copies of $\mathcal{O}(-1)\oplus\mathcal{O}(-1)\to\mathbb{P}^{1}$, with $\alpha_{1,2}\neq 0$. This is called suspended pinched point (SPP) in [OSY11] and [MP25]. The curve from [OSY11] reads as
@@
which matches (1.1) after transforming $e^{x}\to-e^{-x}/Q$ and $e^{y}\to e^{-y}$ with $\alpha_{1}=1/Q,\alpha_{2}=\mu$, for $f=0$. The GW partition function defined as $F_{\geq 2}=\sum_{g\geq 2}\hbar^{2g-2}F_{g}$ has zero radius of convergence in $\hbar$. So, we would need to perform the analysis of [ASTT23] to define various Stokes sectors. In one of those sectors one can recover the partition function of [OSY11]. We have
@@
from which we read off that there are three kinds of stable D2-branes in the type IIA side, with classical (not instanton-corrected) periods $\log Q$, $\log\mu$ and $\log(Q\mu)$ respectively, which can be read off from the associated Picard-Fuchs equations as well [FJ05]. The GV invariants then can be read off as
@@
and all others are zero, where we denote $n^{(g)}_{\beta}$ as the GV invariants, $g$ being the genus and $\beta$ the corresponding curve class in $\mathcal{X}$. It was shown in [BLR21], that exponential networks associated to such geometries compute the GV invariants as $\Omega(\beta+kD0)=n^{(0)}_{\beta}$. In §3, we use the following notation for the 5D indices which for these cases can be read off from the GV invariants, as
@@
where we label the D2 brane charges with the corresponding classical periods and the hyphen denotes the bound states of those D2 branes with several D0’s. Let us also report that the topological string partition function from [OSY11] whose asymptotic expansion in $\hbar$ is (1.6)
@@
In this work, our main result is the following. Using topological recursion which is a universal recursion procedure [EO07b] to compute an infinite family of multidifferentials $\omega_{g,n}$ on $n$ copies of the spectral curve, which actually have their own intrinsic definition that applies to more general classes of curves compared to the mirror curves. We will derive the free energy (2.3) exactly for the geometries in question (“strip geometries”) starting from (1.1). This is strictly a B-model computation. However, by the remodeling conjecture [BKMnP09], this is related to the generating function of the closed GW invariants in the A-model, for the CY 3-fold $\mathcal{Y}$.
Starting from the GW partition function (2.15), we set $q=e^{\hbar}$ to obtain it in the Gopakumar-Vafa form, as was also observed in [Kat06]. Following the computation in [BLR21, Appendix A], one obtains the DT partition function from (2.16) as
@@
Having said this, let us briefly mention the efficacy of topological recursion for the computation of the DT invariants. We will provide more details in the following section, concretely for the geometries at hand. Topological recursion (TR) [EO07b] is a universal procedure to compute an infinite family of multidifferentials $\omega_{g,n}$ indexed by $g,n\in\mathbb{Z}_{>0}$ from some initial data, the spectral curve (1.1). The actual computation of TR is recursive in 2g+n22g+n-2 in the sense that $\omega_{g,n}$ is obtained from all $\omega_{g^{\prime},n^{\prime}}$ with $2g^{\prime}+n^{\prime}-2<2g+n-2$. We are mainly interested here in the free energy denoted by $\omega_{g,0}=F_{g}$ corresponding to the closed string amplitudes. Since TR has this recursive nature, it is almost impossible to obtain closed exact expressions for $\omega_{g,n}$. However, recent developments in TR regarding the so-called xxyy duality [ABDB+22, Hoc24a] provide a new tool for actual computations avoiding the recursive procedure. We will apply this new machinery to derive the closed formulas for the free energy of the mirror curve (1.1). This can be used here efficiently, because the xxyy dual multidifferentials $\omega_{g,n}^{\vee}$ are given by an explicit formula due to $y$ being unramified. In doing so, we have to apply an extension of TR which is called logarithmic topological recursion (Log-TR) [ABDB+24c, Hoc24b]. Thus, we prove that TR derives the expected partition function and/or free energy (closed string amplitudes) for the strip geometries [OSY11]. This sheds light on the interpretation of the xxyy duality in topological recursion in topological string theory and a connection to DT invariants and 5D BPS indices, at least for strip geometries.
@@
As a remark, we would like to point out that by the remodeling conjecture [BKMnP09], $\omega_{g,n}$’s encode the maps from a Riemann surface with genus $g$ and $n$ marked points to $\mathcal{X}$ as $\phi:(\Sigma_{g,n},\partial\Sigma_{g,n})\to(\mathcal{X},L)$. We will come to their computations in a work in preparation [BH], for the case of strip geometries. We will be able to explicitly reproduce the open DT-partition functions [PS19], and we will also propose a quantization program using (Log-)TR and xxyy duality via [HS25].
@@
Finally, we would like to remark that, the class of CY geometries that we analyze in this paper. The strip geometries have been studied using several methods before from different perspectives. In [BLR21] for example, exponential networks were invoked, whereas in [MP25], the computations were done using the 5D BPS quivers which were derived from the B-model side. They were studied using matrix models in [OSY11], or using techniques from crystal melting in [Sul11]. The closed topological string partition function was also derived in [IKP06] by gluing topological vertices. In fact their connection to Nekrasov’s functions [Nek03] were found in [IKP06] and was checked as well in [EK03]. Therefore, there are several results with which we can match our main ingredient that we computed in this paper, the genus $g$ free energy (1.3).
@@
The novelty of this work is that it uses the “remodeling conjecture” of [BKMnP09] to derive explicit free energy expressions for strip geometries. The key technical tool is the application of xxyy duality, which constitutes the main methodological development of this paper. This approach allows us to access the generating functions of Gopakumar–Vafa invariants in a direct way, and thereby compute the associated 5D BPS indices. By comparing with established A-model computations—such as those obtained from 5D BPS quivers or crystal melting models—our B-model derivation provides a concrete test of mirror symmetry. In addition, this method offers an alternative computational framework to exponential networks, and enables nontrivial consistency checks of the results obtained there.
@@
First, one may consider mirror curves from other classes of toric trees as $\mathbb{C}^{3}/(\mathbb{Z}_{m}\times\mathbb{Z}_{n})$ [OSY11]. Those mirror curves will have in TR a nontrivial dual side, that is, all $\omega_{g,n}^{\vee}\neq 0$, but the free energy and the partition function still admit a nice expression [OSY11]. Furthermore, one can consider mirror curves of higher genus, thus having nontrivial 4-cycles. To derive the partition function or the free energy, one has to use the so-called non-perturbative topological recursion, which, however, is not yet properly developed in the context of Log-TR.
Second, one can study the quantum curve arising from TR for strip geometries, which will appear in our upcoming work  [BH]. One should also investigate further geometries with toric diagrams that are not strips. This will provide a direct connection between TR and open DT invariants, following [PS19]. In doing so, it might be possible to understand a correspondence between TR and quivers.
@@
Last but not least, refined GW invariants are well known but not yet obtainable from TR for toric Calabi–Yau threefolds. By formula (2.14), it is actually possible to derive the refined GW invariants for strip geometries from this formula by adding formally $\epsilon_{1}^{2g_{1}-1}$ to $\omega_{g_{1},1}^{\vee}$ and $\epsilon_{2}^{2g_{2}-1}$ to $\omega_{g_{2},1}$. The unrefined case is $\hbar=\epsilon_{1}=-\epsilon_{2}$. Following exactly the computation in Appendix A, we recover refined GW invariants [IKV09]. However, how to modify the original definition of TR to achieve this directly from $\omega_{g,n}$ is not clear to us. It is even not clear if the xxyy duality should hold in a refined setting. This approach to refinement should clearly be distinguished from the one in [KO23].

### v2

@@ Abstract.
Topological string theory partition function gives rise to Gromov-Witten invariants, Donaldson-Thomas invariants and 5D BPS indices. Using the remodeling conjecture, which connects Topological Recursion with topological string theory for toric Calabi–Yau threefolds, we study a more direct connection for the subclass of strip geometries. In doing so, new developments in the theory of topological recursion are applied as its extension to Logarithmic Topological Recursion (Log-TR) and the universal xxyy duality. Through these techniques, our main result in this paper is a direct derivation of all free energies from topological recursion for general strip geometries. In analyzing the expression of free energy, we shed some light on the meaning and the influence of the xxyy duality in topological string theory and its interconnection to GW and DT invariants as well as the 5D BPS index.
@@ 1. Introduction and summary
In this work, our main focus lies on understanding the connection among various kinds of invariants, motivated by enumerative geometry and physics, associated with Calabi-Yau threefolds (CY). Namely, the Gromov-Witten (GW) invariants arising as enumerative invariants in the context of counting stable maps from a Riemann surface $\Sigma_{g}$ with genus $g$ into the CY ($\mathcal{Y}$), in the class $\beta\in H_{2}(\mathcal{Y},\mathbb{Z})$, on the one hand. Through the “remodeling conjecture” [25], (several aspects of which have been firmly established as mathematically proven already [36, 37, 38]) the genus 0 GW invariants can be associated with the B-model geometry of the mirror $\mathcal{X}$, as periods/variation of Hodge-structure, whereas, the higher genera invariants are determined by topological recursion, the main technique that features in this paper. On the other hand, we consider the Donaldson-Thomas (DT) invariants for a CY threefold $\mathcal{Y}$, associated with the counting of coherent sheaves, with stability determined by polarization [56, 21]. A relation between GW invariants and DT invariants is already known due to [64, 65]. As we will see, DT invariants are also closely related to the counting of certain BPS indices in five-dimensional supersymmetric gauge theories, engineered by CY threefolds [72, 73, 15]. A relation between BPS indices and topological recursion was already studied for other geometries in [52, 53]. In the following, we will denote as $\mathcal{X}$ the mirror of the CY threefold $\mathcal{Y}$. In fact, the same counting problem through mirror symmetry can be formulated as the counting of stable special Lagrangian A-branes in the Calabi-Yau threefold ($\mathcal{X}$).
@@
The DT invariants correspond to the counting of BPS states in the 5D $\mathcal{N}=1$ theory obtained by compactifying M-theory on a CY threefold. Starting from M-theory on $\mathcal{Y}\times S^{1}\times\mathbb{R}^{4}$, one arrives at the 5D theory, defined by the geometry of the CY as $T^{\textrm{5D}}[\mathcal{Y}]$ on ${\mathbb{R}}^{4}\times S^{1}$. Now consider M5-branes wrapping $\mathcal{C}_{4}\times S^{1}\times\mathbb{R}$ and M2-branes wrapping $\mathcal{C}_{2}\times{\mathrm{pt}}\times\mathbb{R}$, where $\mathcal{C}_{i}$ are holomorphic submanifolds in $\mathcal{Y}$ (ii corresponds to real dimension) and the second component and the third component indicate the submanifolds that they wrap inside $S^{1}$ and $\mathbb{R}^{4}$. Reducing on $S^{1}$ one obtains a bound state of D4-D2-D0 branes in type IIA string theory corresponding to an object of the derived (bounded) category of coherent sheaves $D^{b}Coh(\mathcal{Y})$ [29, 60, 63, 72]. At the same time in $T^{\textrm{5D}}[\mathcal{Y}]$, the corresponding BPS objects are “monopole strings” wrapping $S^{1}\times\mathbb{R}$ and “instanton particles” wrapping $\mathbb{R}$. This relates the computation of the BPS index for this theory, to the computation of the “rank zero DT invariants” for the Calabi-Yau threefold $\mathcal{Y}$. In this work, for a specific class of geometries, we will provide these 5D BPS indices $\Omega$, starting from the GW partition function.
@@
The type IIA string theory setup for the count of the $\Omega$ is related to the type IIA setup of the Gopakumar-Vafa (GV) through “9-11” flip [29]. Starting from the background $\mathcal{Y}\times S^{1}\times\mathbb{R}^{4}$, one deforms the $\mathbb{R}^{4}$ as a Taub-Nut as $\mathbb{R}^{3}\times S^{1}_{\mathrm{TN}}$. Reducing on the circle $S^{1}_{\mathrm{TN}}$ now, one obtains the relevant type IIA frame where D2 and D0 branes lift to M2 branes and Kaluza-Klein momenta in M-theory. It is possible to include D4 branes which lift to M5 branes. However, for the class of CY threefolds $\mathcal{Y}$ that we will consider, there will be no compact 4-cycle, so we omit their description here.
@@
The GV invariants are intimately related to the GW invariants. This is well-known from both physics and mathematics perspectives [62, 75]. The GW invariants however appear naturally in the context of A-model where one counts stable maps $\phi:\Sigma_{g}\to\mathcal{Y}$. Then the path integral in the A-model localizes to the holomorphic maps which correspond to closed algebraic curves in $\mathcal{Y}$. We will return to their geometric meaning in a short while, after describing the class of the CY threefolds in which our interest lies.
@@
We consider toric CY threefolds $\mathcal{Y}$ which in the A-model side appear as a symplectic quotient $\mathcal{Y}=\mathbb{C}^{k+3}//U(1)^{k}$. The moment map equations give rise to a toric diagram which defines $\mathcal{Y}$. The mirror B-model description was provided in [49], which was $\mathcal{X}=\{A(e^{x},e^{y})=uv\}\subset\mathbb{C}^{4}$. The mirror Calabi-Yau $\mathcal{X}$ then can be considered as a conic fibration on a Riemann surface defined through $A(e^{x},e^{y})=0$. In the particular class of examples that we consider, we focus on the “strip geometries” which belong to the class of “toric trees” [39, 76, 50]. They do not have any compact 4-cycle and the “mirror curve” is given as follows [76]
@@
One way to relate the problem of counting 5D indices $\Omega$ (which are related to the DT invariants) is making the problem slightly more complicated through introducing a toric brane, which we will refer to as Aganagic-Vafa (AV) brane [2, 1, 25]. Consider an M5 brane wrapping $L\times S^{1}\times\mathbb{R}^{2}$, where $L$ corresponds to a special Lagrangian A-brane with topology $\mathbb{C}\times S^{1}$. It was argued in [2] that the complexified moduli space of this toric brane is the mirror curve corresponding to the mirror CY threefold $\mathcal{X}$. This is at the heart of the remodeling conjecture proposed in [25] connecting the B-model computation to topological recursion. Our computations in this paper clarify certain aspects of it, in the context of strip geometries.
However the presence of the AV brane gives rise to a new BPS sector corresponding to a 3D $\mathcal{N}=2$ theory engineered by the codimension two defect coupled to the 5D theory, giving rise to the 3D-5D BPS sector. These were partly understood in [15, 17, 16]. Generalizing the works of [40], it was shown how to compute the DT invariants of $\mathcal{X}$ starting from $(\mathcal{Y},L_{\mathrm{AV}})$ through the procedure of “nonabelianization”. This later led to a geometric definition of DT invariants corresponding to objects in the Fukaya category $D^{b}Fuk(\mathcal{X})$ which are compactly supported special Lagrangian A-branes, [18], providing evidence for a proposal of Joyce [57]. An alternative perspective emerged in [55] and more will be clarified in an upcoming work [14], whereas the computations here will be extended for the case of open GW invariants and connections to open DT invariants will be clarified in [13].
As mentioned above, we focus on toric tree Calabi-Yau (CY) threefolds that do not have compact 4-cycles given by a set of $\mathbb{P}^{1}$’s which have resolutions as either $\mathcal{O}\oplus\mathcal{O}(-2)\to\mathbb{P}^{1}$ or $\mathcal{O}(-1)\oplus\mathcal{O}(-1)\to\mathbb{P}^{1}$. The topological string partition functions for such geometries are known [50, 74]. Let us recall them here:
@@
where we put $s_{i_{k}}=-1$ if the $\mathbb{P}^{1}$ was resolved by $\mathcal{O}(-1,-1)$ and $s_{i_{k}}=+1$ if the resolution is $\mathcal{O}(0,-2)$. Here $\chi$ is the Euler characteristic of the CY and $Q_{i}$ are the complex structure parameters. Looking at the partition function itself, we can write down the toric tree unambiguously, because the topological string partition function can be built up just from the gluing data in this case. Expanding the partition function in terms of the free energy by $Z=\exp\sum_{g}\hbar^{2g-2}F_{g}$ with $q=e^{\hbar}$ yields for $g\geq 2$
@@
for some $k\in\{1,...,\chi-1\}$ depending on the structure of the strip geometry. The 5D BPS invariants depend on the charge $\Gamma(\mathcal{E})={\textrm{ch}}(\mathcal{E})\sqrt{\textrm{Td}(\mathcal{X})}$, for some coherent sheaf $\mathcal{E}$, $\mathcal{X}$ being the CY threefold in the B-model side. From this one can read off the charge vectors $\gamma$ which define the D4-D2-D0 brane-content. Using our proposal, we read off $\Omega(\gamma)$ from (1.3) (in these examples D4-brane charge is zero).
As an example, consider the five punctured sphere obtained from gluing two copies of $\mathcal{O}(-1)\oplus\mathcal{O}(-1)\to\mathbb{P}^{1}$, with $\alpha_{1,2}\neq 0$. This is called suspended pinched point (SPP) in [74] and [68]. The curve from [74] reads as
@@
which matches (1.1) after transforming $e^{x}\to-e^{-x}/Q$ and $e^{y}\to e^{-y}$ with $\alpha_{1}=1/Q,\alpha_{2}=\mu$, for $f=0$. The GW partition function defined as $F_{\geq 2}=\sum_{g\geq 2}\hbar^{2g-2}F_{g}$ has zero radius of convergence in $\hbar$. So, we would need to perform the analysis of [11] to define various Stokes sectors. In one of those sectors one can recover the partition function of [74]. We have
@@
from which we read off that there are three kinds of stable D2-branes in the type IIA side, with classical (not instanton-corrected) periods $\log Q$, $\log\mu$ and $\log(Q\mu)$ respectively, which can be computed from the associated Picard-Fuchs equations as well [39]. The GV invariants then can be read off as
@@
and all others are zero, where we denote $n^{(g)}_{\beta}$ as the GV invariants, $g$ being the genus and $\beta$ the corresponding curve class in $\mathcal{X}$. It was shown in [17], that exponential networks associated to such geometries compute the GV invariants as $\Omega(\beta+kD0)=n^{(0)}_{\beta}$. In §3, we use the following notation for the 5D indices which for these cases can be read off from the GV invariants, as
@@
where we label the D2 brane charges with the corresponding classical periods and the hyphen denotes the bound states of those D2 branes with several D0’s. Let us also report that the topological string partition function from [74] whose asymptotic expansion in $\hbar$ is (1.6)
@@
In this work, our main result is the following. We use topological recursion which is a universal recursion procedure [35] to compute an infinite family of multidifferentials $\omega_{g,n}$ on $n$ copies of the spectral curve, which actually have their own intrinsic definition that applies to more general classes of curves compared to the mirror curves. We will derive the free energy (2.3) exactly for the geometries in question (“strip geometries”) starting from (1.1). This is strictly a B-model computation. However, by the remodeling conjecture [25], this is related to the generating function of the closed GW invariants in the A-model, for the CY threefold $\mathcal{Y}$.
Starting from the GW partition function (2.15), we set $q=e^{\hbar}$ to obtain it in the Gopakumar-Vafa form, as was also observed in [58]. Following the computation in [17, Appendix A], one obtains the DT partition function from (2.16) as
@@
Having said this, let us briefly mention the efficacy of topological recursion for the computation of the DT invariants. We will provide more details in the following section, concretely for the geometries at hand. As stated above, topological recursion (TR) [35] provides a universal framework to associate an infinite family of multidifferentials $\omega_{g,n}$ indexed by $g,n\in\mathbb{Z}_{>0}$ to certain initial data, the spectral curve (1.1). The actual computation of TR is recursive in 2g+n22g+n-2 in the sense that $\omega_{g,n}$ is obtained from all $\omega_{g^{\prime},n^{\prime}}$ with $2g^{\prime}+n^{\prime}-2<2g+n-2$. We are mainly interested here in the free energy denoted by $\omega_{g,0}=F_{g}$ corresponding to the closed string amplitudes. Since TR has this recursive nature, it is almost impossible to obtain closed exact expressions for $\omega_{g,n}$. However, recent developments in TR regarding the so-called xxyy duality [3, 46] provide a new tool for actual computations avoiding the recursive procedure. We will apply this new machinery to derive the closed formulas for the free energy of the mirror curve (1.1). This can be implemented here efficiently, because the xxyy dual multidifferentials $\omega_{g,n}^{\vee}$ are given by an explicit formula due to $y$ being unramified. In doing so, we have to apply an extension of TR which is called logarithmic topological recursion (Log-TR) [6, 47]. Thus, we demonstrate that for strip geometries, using TR one can derive the GW partition functions from which one can recover the corresponding DT partiton functions, matching with for example [74]. This sheds light on the interpretation of the xxyy duality in topological recursion in the context of topological string theory and a connection to DT invariants and 5D BPS indices, at least for strip geometries.
@@
As a remark, we would like to point out that by the remodeling conjecture [25], $\omega_{g,n}$’s encode the maps from a Riemann surface with genus $g$ and $n$ marked points to $\mathcal{X}$ as $\phi:(\Sigma_{g,n},\partial\Sigma_{g,n})\to(\mathcal{X},L)$. We will come to their computations in a work in preparation [13], for the case of strip geometries. We will be able to explicitly reproduce the open DT-partition functions [76], and we will also propose a quantization program using (Log-)TR and xxyy duality via [43].
@@
Finally, we would like to mention the class of CY geometries analyzed in this paper. The strip geometries have been studied using several methods before from different perspectives. In [17] for example, exponential networks were invoked, whereas in [68], the computations were done using the 5D BPS quivers which were derived from the B-model side. They were studied using matrix models in [74], or using techniques from crystal melting in [77]. The closed topological string partition function was also derived in [50] by gluing topological vertices. In fact their connection to Nekrasov’s functions [70] were found in [50] and was checked as well in [31]. Therefore, there are several results with which we can match our main ingredient that we computed in this paper, the genus $g$ free energy (1.3).
@@
The novelty of this work is that it uses the “remodeling conjecture” of [25] to derive explicit free energy expressions for strip geometries. The key technical tool is the application of xxyy duality, which constitutes the main methodological ingredient of this paper. This approach allows us to access the generating functions of Gopakumar–Vafa invariants in a direct way, and thereby compute the associated 5D BPS indices. By comparing with established A-model computations—such as those obtained from 5D BPS quivers or crystal melting models—our B-model derivation provides a concrete test of mirror symmetry. In addition, this method offers an alternative computational framework to exponential networks, and enables nontrivial consistency checks of the results obtained there.
@@
First, one may consider mirror curves from other classes of toric trees as $\mathbb{C}^{3}/(\mathbb{Z}_{m}\times\mathbb{Z}_{n})$ [74]. Those mirror curves will have in TR a nontrivial dual side, that is, all $\omega_{g,n}^{\vee}\neq 0$, but the free energy and the partition function still admit a nice expression [74]. Furthermore, one can consider mirror curves of higher genus, thus having nontrivial 4-cycles. To derive the partition function or the free energy, one has to use the so-called non-perturbative topological recursion, which, however, is not yet properly developed in the context of Log-TR.
Second, one can study the quantum curve arising from TR for strip geometries, which will appear in our upcoming work  [13]. One should also investigate further geometries with toric diagrams that are not strips. This will provide a direct connection between TR and open DT invariants, following [76]. In doing so, it might be possible to understand a correspondence between TR and quivers.
@@
Last but not least, refined GW invariants are well known but not yet obtainable from TR for toric Calabi–Yau threefolds. By formula (2.14), it is actually possible to derive the refined GW invariants for strip geometries from this formula by adding formally $\epsilon_{1}^{2g_{1}-1}$ to $\omega_{g_{1},1}^{\vee}$ and $\epsilon_{2}^{2g_{2}-1}$ to $\omega_{g_{2},1}$. The unrefined case is $\hbar=\epsilon_{1}=-\epsilon_{2}$. Following exactly the computation in Appendix A, we recover refined GW invariants [51]. However, how to modify the original definition of TR to achieve this directly from $\omega_{g,n}$ is not clear to us. It is even not clear if the xxyy duality should hold in a refined setting. This approach to refinement should clearly be distinguished from the one in [59].